# Tales of stellar and binary co-evolution, told by stellar oscillations and age estimates in main-sequence and red-giant stars

### v1

@@ 5 Evolution of the radial velocities
From the evolution of with respect to , Badenes et al. (2018, Fig. 5) reported a significant attrition of the peak-to-peak amplitude from 200 km/s on the MS to 30 km/s at the RGB tip. To test if we find a similar accumulated effect in the seismically characterized APOKASC 3 and APO-K2 samples, we use provided in the TBO (Gaia Collaboration et al., 2023a).
@@ 5.1 RV vs orbital period and stellar radius
Figure 8 (top panel, see also Fig. 6) depicts an apparent reduction of the semi-RV amplitude from the MS is where semi-amplitudes of nearly 100 km/s are found. On the low-luminous RGB, K 30 km/s are found, but are reduced to K 20 km/s on the luminous RGB. Most RC and 2RC systems are found with K 10 km/s, with a few systems reaching 20 km/s.
When comparing these numbers with Badenes et al. (2018), we need to take into account that they are using the peak-to-peak variation, which is up to twice the semi-amplitude K. Therefore, the values presented in Fig. 8 are in good agreement with large-sample spectroscopic studies. The detection threshold value for a significant RV variation with the Gaia RVS instrument is 0.3 and 1.4 km/s for bright and faint stars (Katz et al., 2023). Given the typical in red-giant binary systems and the brightness of targets in the Kepler space photometry, this RV floor is sufficient. Similarly, a reduction of is also seen as a function of the stellar radius (Fig. 8). Figure 6 shows that we find a cutoff of the oscillating sample of RGB stars at a fractional radius of 20%. We further note that MS+SG and the core-He burning (RC+2RC) stars are well below this threshold (Fig. 6).
@@
To test if red-giant primaries are close to filling their Roche lobes, we estimated the critical period, PRL. We follow the formalism of Eggleton (1983),
where is a function of the mass ratio that gives the Roche lobe radius relative to the orbital separation, and is for the mass ratio, which we assume to be at unity. Following Eggleton (1983), can be approximated by 0.38 for the unity mass ratio. Miglio et al. (2014) pointed out that within a few percent deviation, 1 is expected for the majority of red-giant binary stars. Furthermore, Eggleton (1983) noted that for 0.01 q 1, the function varies by less than 3%. The resulting velocities for giants with 0.5, 1, and 2 M⊙, would be filling their Roche lobes are calculated from Eq. 2 and Eq. 3, assuming an edge-on inclination ( = 90) and circular orbits. The gap between the detected systems and the Roche-lobe limit can be attributed to the suppression of modes due to enhanced activity driven by the increasing strength of the equilibrium tide. Therefore, those systems are far from filling their Roche lobe.
5.2 RV vs stellar mass
@@ 6 Discussions and conclusions
In this paper, we investigated the binary properties of red-giant stars by combining asteroseismic constraints from the APOKASC 3 and APO-K2 catalogs with binary solutions from Gaia DR3’s NSS catalog. The asteroseismically inferred stellar masses and evolutionary states provide key diagnostics for characterizing binary systems hosting solar-like oscillators. This cross-matched dataset enabled a comprehensive analysis of how binary fraction, stellar mass distribution, and orbital parameters evolve across different evolutionary stages, offering a holistic view of stellar–binary co-evolution.
Our results demonstrate a significant reduction in the binary fraction as stars ascend the RGB and transition into the core-helium-burning phase. To account for distinct evolutionary pathways, we divided the sample based on asteroseismic masses into low-mass stars (M 1.8 M⊙), which develop fully degenerate helium cores, and intermediate-mass stars (M 1.8 M⊙), which ignite helium under non-degenerate conditions. We find that approximately 31.5% and 40.7% of solar-like stars on the main sequence, with detected and non-detected solar-like oscillations, respectively, are members of binary systems. This is a noteworthy result, as magnetic activity, often enhanced in close binaries, has been suggested to reduce solar-like oscillations in dwarfs. Using as a proxy for stellar luminosity, we further subdivided the RGB into low- and high-luminosity regimes. Our analysis reveals a significant reduction of 69% in the binary fraction from the main sequence to the low-luminosity RGB, followed by a further 38% decrease from the low- to high-luminosity RGB. Interestingly, no significant change in the binary fraction is found between the high-luminosity RGB and RC, suggesting that most binary attrition or interaction occurs earlier during the RGB ascent. For the intermediate-mass stars (M 1.8 M⊙), we find a substantially higher binary fraction. This is consistent with the intrinsically higher binary occurrence rate (approx. 8 to 14%), observed among early F- and A-type stars on the main sequence. However, due to the smaller sample size in this mass regime, we are currently unable to robustly test for evolutionary changes in binary fractions across the MS.
@@
The distribution of orbital parameters reveals a clear evolutionary trend. While main-sequence stars span a wide range of orbital periods, resembling the expected initial distribution, systems hosting RGB and core-helium-burning stars are predominantly found at longer periods and with reduced eccentricities. RC stars are typically found in systems with orbital periods longer than 500 days. In contrast, more massive 2RC are found at shorter periods, down to 200 days. This reduction is interpreted as evidence of a binary interaction or disruption and is consistent with the fact that intermediate-mass stars ignite He at smaller radii on the RGB, allowing tighter systems to avoid interaction.
@@
As orbital periods approach 800 to 1,000 days, corresponding to semi-major axes comparable to red giant radii, binary co-evolution becomes increasingly complex. In this regime, the otherwise simplifying assumption of co-evolution of both stellar components and the binary orbit is challenged. In particular, age estimates from single-star models may no longer be reliable for post-interaction products.
Our results provide clear observational evidence for mass- and evolution-dependent attrition of binary systems during the red-giant phase. These findings have important implications for binary population synthesis and the interpretation of chemical and rotational anomalies in evolved stars. In particular, the ages of helium-core burning stars may be misleading if prior mass loss or transfer is unaccounted for. Red-clump systems that have undergone substantial interaction are therefore unsuitable for calibrating stellar evolution models based on single-star physics. This study underscores the powerful synergy between asteroseismology and Gaia in unraveling stellar and orbital co-evolution, by providing valuable insights for more refined binary population synthesis and age-dating techniques.
The size of available samples of oscillating red giants in binaries is expected to grow substantially with upcoming data releases and forthcoming space missions. The extended time baseline and epoch spectroscopy provided by Gaia DR4 will significantly improve the census and precision of astrometric and spectroscopic binary solutions, particularly for long-period systems. The Science Calibration and Validation Plato Input Catalog (scvPIC) of ESA’s PLATO mission (PLAnetary Transits and Oscillations of stars; Rauer et al., 2025) will further increase the number of oscillating stars in eclipsing and astrometric binary systems identified by Gaia. Additional contributions will come from NASA’s Roman Space Telescope (Huber et al., 2023) and the Chinese Earth 2.0 mission (Ge et al., 2022), which will extend the ensemble of solar-like oscillators across diverse regions of the Milky Way. Together, these missions will provide the necessary data to probe binary evolution in the context of galactic archaeology. This underscores the need for a holistic picture of stellar and binary co-evolution. This is essential for accurately determining stellar ages and unlocking the full potential of red giants as tracers of Galactic history.
@@ Acknowledgements.
The author thanks the referee for their valuable comments, which helped improve the article. The author thanks the people behind the ESA Gaia, NASA Kepler, NASA TESS, and APOGEE projects. PGB thanks Santi Cassisi, Thomas Masseron, Joel Zinn, Savita Mathur, Diego Godoy Rivera, and Jaroslav Merc, for fruitful discussions on the catalogs and the paper manuscript. PGB acknowledges support by the Spanish Ministry of Science and Innovation with the Ramón y Cajal fellowship number RYC-2021-033137-I and the number MRR4032204, as well as the proyecto plan nacional PLAtoSOnG (grant no. PID2023-146453NB-100, PI: Beck). This project was supported by “UNDARK” of the Widening participation and spreading excellence programme (project number 101159929, PI: Camalich).This work has made use of data from the European Space Agency (ESA) mission Gaia, processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Data collected with the Kepler mission, obtained from the MAST data archive at the Space Telescope Science Institute (STScI) was used. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. Software: Python (Van Rossum & Drake, 2009), numpy (Oliphant, 2006; Harris et al., 2020), matplotlib (Hunter, 2007), scipy (Virtanen et al., 2020), pandas (pandas development team, 2020; Wes McKinney, 2010), Astroquery (Ginsburg et al., 2019).
@@ A.1 Notes on CDS Table: Gaia DR3 binary systems in the APOKASC3
nu_max: Frequency of maximum oscillation power (Hz), derived from Kepler photometry. -
nu_max_err: Uncertainty in the frequency of maximum oscillation power (Hz). -
Delta_nu: Large frequency separation between consecutive overtone modes (Hz). -
Delta_nu_err: Uncertainty in the large frequency separation (Hz). -
@@
K1: Radial-velocity semi-amplitude (km s-1) of the primary component derived from the Gaia orbital solution. -
RUWE: Renormalised Unit Weight Error from the Gaia DR3 astrometric solution; indicator of astrometric excess noise. -
NLAC:[0,1]Flag identifying if a source is listed (Flag=1) or not-listed (Flag=0) with non-linear or acceleration solutions in the Gaia NSS catalog. -
@@ A.2 Notes on CDS Table: Gaia DR3 binary systems in the APO-K2
nu_max: Frequency of maximum oscillation power (Hz), derived from K2 photometry (Zinn et al. 2022). -
nu_max_err: Uncertainty in the frequency of maximum oscillation power (Hz). -
Delta_nu: Large frequency separation between consecutive overtone modes (Hz), derived from asteroseismic analysis of K2 light curves. -
Delta_nu_err: Uncertainty in the large frequency separation (Hz). -
@@
K1: Radial-velocity semi-amplitude (km s-1) of the primary component derived from the Gaia orbital solution. -
@@
RUWE: Renormalised Unit Weight Error from the Gaia DR3 astrometric solution; indicator of astrometric excess noise. -
NLAC:[0,1]Flag indicating whether a source is listed (1) or not listed (0) with non-linear or acceleration solutions in the Gaia NSS catalog.

### v2

@@ 5 Evolution of the radial velocities
From the evolution of with respect to , Badenes et al. (2018, Fig. 5) reported a significant attrition of the peak-to-peak amplitude from 200 km/s on the MS to 30 km/s at the RGB tip. To test if we find a similar accumulated effect in the seismically characterized APOKASC 3 and APO-K2 samples, we used provided in the TBO (Gaia Collaboration et al., 2023a).
@@ 5.1 RV vs orbital period and stellar radius
Figure 8 (top panel, see also Fig. 6) depicts an apparent reduction in the semi-RV amplitude from the MS is found. On the low-luminous RGB, K 30 km/s are found, but are reduced to K 20 km/s on the luminous RGB. Most RC and 2RC systems are found with K 10 km/s, with a few systems reaching 20 km/s.
When comparing these numbers with Badenes et al. (2018), we need to take into account that they are using the peak-to-peak variation, which is up to twice the semi-amplitude, K. Therefore, the values presented in Fig. 8 are in good agreement with large-sample spectroscopic studies. The detection threshold value for a significant RV variation with the Gaia RVS instrument is 0.3 and 1.4 km/s for bright and faint stars (Katz et al., 2023). Given the typical in red-giant binary systems and the brightness of targets in the Kepler space photometry, this RV floor is sufficient. Similarly, a reduction of is also seen as a function of the stellar radius (Fig. 8). Figure 6 shows that we find a cutoff of the oscillating sample of RGB stars at a fractional radius of 20%. We further note that MS+SG and the core-He burning (RC+2RC) stars are well below this threshold (Fig. 6).
@@
To test if red-giant primaries are close to filling their Roche lobes, we estimated the critical period, PRL. We followed the formalism of Eggleton (1983),
where is a function of the mass ratio that gives the Roche lobe radius relative to the orbital separation, and is for the mass ratio, which we assume to be at unity. Following Eggleton (1983), can be approximated by 0.38 for the unity mass ratio. Miglio et al. (2014) pointed out that within a deviation of a few percent, 1 is expected for the majority of red-giant binary stars. Furthermore, Eggleton (1983) noted that for 0.01 q 1, the function varies by less than 3%. The resulting velocities for giants with 0.5, 1, and 2 M⊙, would be filling their Roche lobes are calculated from Eq. 2 and Eq. 3, assuming an edge-on inclination ( = 90) and circular orbits. The gap between the detected systems and the Roche-lobe limit can be attributed to the suppression of modes due to enhanced activity driven by the increasing strength of the equilibrium tide. Therefore, those systems are far from filling their Roche lobe.
5.2 RV versus stellar mass
@@ 6 Discussions and conclusions
In this paper, we investigated the binary properties of red-giant stars by combining asteroseismic constraints from the APOKASC 3 and APO-K2 catalogs with binary solutions from Gaia DR3’s NSS catalog. The asteroseismically inferred stellar masses and evolutionary states provide key diagnostics for characterizing binary systems hosting solar-like oscillators. This cross-matched dataset enabled a comprehensive analysis of how binary fraction, stellar mass distribution, and orbital parameters evolve across different evolutionary stages, offering a holistic view of stellar–binary coevolution.
Our results demonstrate a significant reduction in the binary fraction as stars ascend the RGB and transition into the core-helium-burning phase. To account for distinct evolutionary pathways, we divided the sample based on asteroseismic masses into low-mass stars (M 1.8 M⊙), which develop fully degenerate helium cores, and intermediate-mass stars (M 1.8 M⊙), which ignite helium under non-degenerate conditions. We find that approximately 31.5% and 40.7% of solar-like stars on the MS, with detected and non-detected solar-like oscillations, respectively, are members of binary systems. This is a noteworthy result, as magnetic activity, often enhanced in close binaries, has been suggested to reduce solar-like oscillations in dwarfs. Using as a proxy for stellar luminosity, we further subdivided the RGB into low- and high-luminosity regimes. Our analysis reveals a significant reduction of 69% in the binary fraction from the MS to the low-luminosity RGB, followed by a further 38% decrease from the low- to high-luminosity RGB. Interestingly, no significant change in the binary fraction is found between the high-luminosity RGB and RC, suggesting that most binary attrition or interaction occurs earlier during the RGB ascent. For the intermediate-mass stars (M 1.8 M⊙), we find a substantially higher binary fraction. This is consistent with the intrinsically higher binary occurrence rate (approx. 8 to 14%), observed among early F- and A-type stars on the MS. However, due to the smaller sample size in this mass regime, we are currently unable to robustly test for evolutionary changes in binary fractions across the MS.
@@
The distribution of orbital parameters reveals a clear evolutionary trend. While MS stars span a wide range of orbital periods, resembling the expected initial distribution, systems hosting RGB and core-helium-burning stars are predominantly found at longer periods and with reduced eccentricities. RC stars are typically found in systems with orbital periods longer than 500 days. In contrast, more massive 2RCs are found at shorter periods, down to 200 days. This reduction is interpreted as evidence of a binary interaction or disruption and is consistent with the fact that intermediate-mass stars ignite He at smaller radii on the RGB, allowing tighter systems to avoid interaction.
@@
As orbital periods approach 800 to 1,000 days, corresponding to semimajor axes comparable to red-giant radii, binary coevolution becomes increasingly complex. In this regime, the otherwise simplifying assumption of coevolution of both stellar components and the binary orbit is challenged. In particular, age estimates from single-star models may no longer be reliable for post-interaction products.
Our results provide clear observational evidence of mass- and evolution-dependent attrition of binary systems during the red-giant phase. These findings have important implications for binary population synthesis and the interpretation of chemical and rotational anomalies in evolved stars. In particular, the ages of helium-core burning stars may be misleading if prior mass loss or transfer is unaccounted for. Red-clump systems that have undergone substantial interaction are therefore unsuitable for calibrating stellar evolution models based on single-star physics. This study underscores the powerful synergy between asteroseismology and Gaia in unraveling stellar and orbital coevolution, by providing valuable insights for more refined binary population synthesis and age-dating techniques.
The size of available samples of oscillating red giants in binaries is expected to grow substantially with upcoming data releases and forthcoming space missions. The extended time baseline and epoch spectroscopy provided by Gaia DR4 will significantly improve the census and precision of astrometric and spectroscopic binary solutions, particularly for long-period systems. The Science Calibration and Validation Plato Input Catalog (scvPIC) of ESA’s PLATO mission (PLAnetary Transits and Oscillations of stars; Rauer et al., 2025) will further increase the number of oscillating stars in eclipsing and astrometric binary systems identified by Gaia. Additional contributions will come from NASA’s Roman Space Telescope (Huber et al., 2023) and the Chinese Earth 2.0 mission (Ge et al., 2022), which will extend the ensemble of solar-like oscillators across diverse regions of the Milky Way. Together, these missions will provide the necessary data to probe binary evolution in the context of galactic archaeology. This underscores the need for a holistic picture of stellar and binary coevolution. This is essential for accurately determining stellar ages and unlocking the full potential of red giants as tracers of Galactic history.
@@ Acknowledgements.
The author thanks the referee for their valuable comments, which helped improve the article. The author thanks the people behind the ESA Gaia, NASA Kepler, NASA TESS, and APOGEE projects. PGB thanks Marc Pinsonneault, Santi Cassisi, Thomas Masseron, Joel Zinn, Savita Mathur, Diego Godoy Rivera, and Jaroslav Merc, for fruitful discussions on the catalogs and the paper manuscript. PGB acknowledges support by the Spanish Ministry of Science and Innovation with the Ramón y Cajal fellowship number RYC-2021-033137-I and the number MRR4032204, as well as the proyecto plan nacional PLAtoSOnG (grant no. PID2023-146453NB-100, PI: Beck). This project was supported by “UNDARK” of the Widening participation and spreading excellence programme (project number 101159929, PI: Camalich). This work was supported by the International Space Science Institute (ISSI) in Bern (project 24-629).This work has made use of data from the European Space Agency (ESA) mission Gaia, processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Data collected with the Kepler mission, obtained from the MAST data archive at the Space Telescope Science Institute (STScI) was used. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. Software: Python (Van Rossum & Drake, 2009), numpy (Oliphant, 2006; Harris et al., 2020), matplotlib (Hunter, 2007), scipy (Virtanen et al., 2020), pandas (pandas development team, 2020; Wes McKinney, 2010), Astroquery (Ginsburg et al., 2019).
@@ A.1 Notes on CDS Table: Gaia DR3 binary systems in the APOKASC3
nu_max: Frequency of maximum oscillation power (µHz), derived from Kepler photometry. -
nu_max_err: Uncertainty in the frequency of maximum oscillation power (µHz). -
Delta_nu: Large frequency separation between consecutive overtone modes (µHz). -
Delta_nu_err: Uncertainty in the large frequency separation (µHz). -
@@
K1: RV semi-amplitude (km s-1) of the primary component derived from the Gaia orbital solution. -
K1_err: Uncertainty in the RV semi-amplitude (km s-1). -
RUWE: Renormalized unit weight error from the Gaia DR3 astrometric solution; indicator of astrometric excess noise. -
NLAC:[0,1]Flag identifying if a source is listed (Flag=1) or not-listed (Flag=0) with nonlinear or acceleration solutions in the Gaia NSS catalog. -
@@ A.2 Notes on CDS Table: Gaia DR3 binary systems in the APO-K2
nu_max: Frequency of maximum oscillation power (µHz), derived from K2 photometry (Zinn et al. 2022). -
nu_max_err: Uncertainty in the frequency of maximum oscillation power (µHz). -
Delta_nu: Large frequency separation between consecutive overtone modes (µHz), derived from asteroseismic analysis of K2 light curves. -
Delta_nu_err: Uncertainty in the large frequency separation (µHz). -
@@
K1: RV semi-amplitude (km s-1) of the primary component derived from the Gaia orbital solution. -
@@
RUWE: Renormalized unit weight error from the Gaia DR3 astrometric solution; indicator of astrometric excess noise. -
NLAC:[0,1]Flag indicating whether a source is listed (1) or not listed (0) with nonlinear or acceleration solutions in the Gaia NSS catalog.